Required information

[The following information applies to the questions displayed below.]

Tyrell Co. entered into the following transactions involving short-term liabilities.

Year 1

Apr. 20 Purchased $38,500 of merchandise on credit from Locust, terms n/30.
May 19 Replaced the April 20 account payable to Locust with a 90-day, 7%, $35,000 note payable along with paying $3,500 in cash.
July 8 Borrowed $69,000 cash from NBR Bank by signing a 120-day, 11%, $69,000 note payable.
__?__ Paid the amount due on the note to Locust at the maturity date.
__?__ Paid the amount due on the note to NBR Bank at the maturity date.
Nov. 28 Borrowed $33,000 cash from Fargo Bank by signing a 60-day, 8%, $33,000 note payable.
Dec. 31 Recorded an adjusting entry for accrued interest on the note to Fargo Bank.


Year 2

__?__ Paid the amount due on the note to Fargo Bank at the maturity date.Chapter 9 Problems i Saved On January 8, the end of the first weekly pay period of the year, Regis Companys employees earned
Answer #1

one question allowed per post , please post the second question in separate post

1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 7% * 90/360 = 613
NBR 69,000 * 11% * 120/360 = 2530
Fargo 33,000 * 8% * 60/360 = 440
3) Amount in adjusting entry
42,000*8%*33/360
242
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 33,000 * 8% * 33/360 = 242
4) interest expense to be recorded in 2017
198
principal * Rate * time = interest
interest to recorded in 2018 33,000 * 8% * 27/360 = 198
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 38,500
Accounts payable 38,500
19-May Accounts payable 38,500
cash 3,500
notes payable 35,000
8-Jul Cash 69,000
notes payable 69,000
17-Aug notes payable 35,000
interest expense 613
cash 35,613
5-Nov notes payable 69,000
interest expense 2,530
cash 71,530
28-Nov Cash 33,000
notes payable 33,000
31-Dec interest expense 242
interest payable 242
2017
27-Jan notes payable 33,000
interest payable 242
interest expense 198
cash 33,440